OpenStudy (anonymous):

You and your friends decide to go bowling on a Friday evening, it costs $2.25 to rent a pair of shoes and $2.75 to bowl one game. the function is represented by f(x) = 2.75x + 2.25 determine the total cost if you bowled 7 games

OpenStudy (anonymous):

would it be $21.50 ?

OpenStudy (anonymous):

Given: \[\huge f(x) =2.75x +2.25\] plugging x=7 in the above function we find; \[\huge f(7) =2.75 \times 7 +2.25 \] \[\huge f(7)=19.25+2.25 =21.50\] @trever74 So you are right..

OpenStudy (anonymous):

thanks so much